1. . A man walks 30.0 miles eastward and then 50.0 miles northward. Solve for the magnitude and direction of the resultant displacement by analytical method/ Pythagorean theorem.
1
Expert's answer
2021-07-27T02:00:04-0400

The magnitude of the resultant displacement

"d=\sqrt{30.0^2+50.0^2}=58.3\:\rm miles"

The direction of the resultant displacement

"\theta=\tan^{-1}\frac{50.0}{30.0}=59.0^{\circ}\;\rm N\;of\; E"